# GEM: Gaussian Embedding Modeling for Out-of-Distribution Detection in GUI Agents

### v3

@@ Abstract
Graphical user interface (GUI) agents have recently emerged as an intriguing paradigm for human-computer interaction, capable of automatically executing user instructions to operate intelligent terminal devices. However, when encountering out-of-distribution (OOD) instructions that violate environmental constraints or exceed the current capabilities of agents, GUI agents may suffer task breakdowns or even pose security threats. Therefore, effective OOD detection for GUI agents is essential. Traditional OOD detection methods perform suboptimally in this domain due to the complex embedding space and evolving GUI environments. In this work, we observe that the in-distribution input semantic space of GUI agents exhibits a clustering pattern with respect to the distance from the centroid. Based on the finding, we propose GEM, a novel method based on fitting a Gaussian mixture model over input embedding distances extracted from the GUI agent that reflect its capability boundary. Evaluated on eight datasets spanning smartphones, computers, and web browsers, our method achieves an average accuracy improvement of 23.70% over the best-performing baseline while only increasing training time by 4.9% and testing time by 6.5%. We also experimentally demonstrate that GEM can improve the step-wise success rate by 9.40% by requesting assistance from the cloud model when encountering OOD samples. Analysis verifies the generalization ability of our method through experiments on nine different backbones. The codes are available at https://github.com/Wuzheng02/GEM-OODforGUIagents.
@@ Introduction
Recently, graphical user interface (GUI) agents (Zhang et al. 2024a; Liu et al. 2025b; Shi et al. 2025) have emerged as an intriguing paradigm to human-computer interaction, capable of autonomously executing user instructions and performing human-like control on intelligent terminal devices such as smartphones, computers, and web browsers. The common approach to building GUI agents involves post-training multimodal large language models (MLLMs) using high-quality trajectory data to enhance key task capabilities such as perception (Li et al. 2023; Wang et al. 2023b), reasoning (Yao et al. 2023; Wei et al. 2022), and reflection (Liu et al. 2025c; Hu et al. 2025). Despite notable advances in instruction following, GUI agents remain vulnerable to out-of-distribution (OOD) risks—executing instructions that violate environmental constraints (e.g., non-existent functions or applications) or exceed the agent’s current capabilities. These failures can lead to task breakdowns or even pose security threats. In real-world applications, OOD risks for GUI agents come with two primary forms: (i) Internalization-OOD, where the agent operates in domain-specific environments (e.g., a particular type of smartphones or vehicle cabins) but incorrectly assumes the presence of unsupported capabilities; (ii) Extrapolation-OOD, where the agent encounters instructions tied to dynamic or evolving environments (e.g., new third-party applications).
As illustrated in Figure 1, following OOD instructions and executing an incorrect action path—such as unintentionally resetting a smartphone—can result in critical failures. Therefore, it is essential for GUI agents to incorporate OOD detection mechanisms that identify tasks beyond their supported scope. This not only mitigates potential operational risks but also enables targeted enhancements to agent capabilities through continued research and development.
Popular OOD detection methods can be broadly categorized (Malinin and Gales 2018) into two types: embedding-based approaches (Lee et al. 2018) and model uncertainty-based approaches (Hendrycks and Gimpel 2016). These methods have demonstrated effectiveness in traditional (M)LLM tasks such as summarization (Ren et al. 2022), visual question answering (Kervadec et al. 2021), mathematical reasoning (Wang et al. 2024b), and text generation (Wu et al. 2022). Directly applying these existing OOD detection techniques to the GUI agent domain results in suboptimal performance. Notably, even with the classification threshold calibrated to correctly identify 95% of in-distribution (ID) tasks, the best-performing method still misclassifies between 31.75% and 48.92% of OOD tasks as ID.
@@
(i) Complex Embedding Space: The inputs to GUI agents are inherently complex (Ma, Zhang, and Zhao 2024). Concretely, GUI screens typically contain densely populated UI components (Zhang et al. 2024a), leading to higher information density than traditional MLLM tasks. Besides, the diversity of user instructions further complicates the inference of user intent. This complexity substantially increases the difficulty of effective OOD detection for GUI agents.
(ii) Evolving Environments: GUI environments frequently change due to system upgrades or the installation of new third-party applications. As a result, GUI agents must continually adapt their capabilities (Wang et al. 2025; Liu et al. 2025a). This evolving nature complicates OOD detection by demanding both accurate capability assessment and temporal adaptability. Consequently, reliance on static external classifiers becomes increasingly impractical.
To address the challenges above, we propose GEM, a novel method based on fitting a Gaussian mixture model (GMM) (Reynolds et al. 2009) over input embedding distances extracted from the GUI Agent that reflect its capability boundary. Specifically, we first extract the input embeddings from the encoder layer of the GUI agent on its training data, and fit them into a high-dimensional hypersphere. We then compute the L2-norm distances of all embeddings relative to the centroid of the hypersphere. Under bayesian information criterion (BIC) (Neath and Cavanaugh 2012) supervision, we fit a GMM to these distances and define the OOD detection boundary as a configurable number of standard deviations away from each GMM cluster center. Experiments across eight GUI agent datasets spanning smartphones, computers, and web browsers platforms show that our method consistently outperforms all traditional OOD detection methods. And we provide a possible method to use GEM to improve the end-to-end results for GUI agents.
@@
Smartphone Computer Web browser
Method AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$
@@
With the advancement of MLLMs (Liu et al. 2023; Achiam et al. 2023), numerous GUI agent foundation models (Hong et al. 2024; Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) have emerged, capable of executing user instructions across smartphones, computers, and web browsers. For the technical framework, prompt-based approaches (Wang et al. 2025; Jiang et al. 2025) have been developed, leveraging closed-source models to construct agent systems that fulfill user instructions. Additionally, researchers have also explored pre-training (Ma, Zhang, and Zhao 2024; Wu et al. 2025b), supervised fine-tuning (SFT) (Zhang and Zhang 2024; Wu et al. 2025a), and reinforcement learning (RL) (Zhou et al. 2024; Liu et al. 2024; Wang et al. 2024a; Xia and Luo 2025) techniques to further enhance the ability of GUI agents to complete user instructions. Regardless of their advancements, GUI agents inevitably face capability limitations. In complex and dynamic real-world environments (Cheng et al. 2025a; Guo et al. 2025), GUI agents are prone to encountering OOD situations.
@@
For OOD detection in MLLMs, it is necessary to jointly consider both visual and textual modalities, in contrast to traditional computer vision OOD detection, which only requires modeling the visual modality (Liu et al. 2020; Ren et al. 2019), or LLMs OOD detection, which focuses on textual modality (Wang et al. 2024b; Ren et al. 2022). In the context of MLLMs, some researchers have proposed using maximum concept matching (Ming et al. 2022) to characterize OOD uncertainty, while others have approached the problem by training models on ID datasets and classifying whether an input image belongs to an unknown category (Wang et al. 2023a). However, OOD detection for MLLMs remains a challenging research area (Dong et al. 2024; Lu et al. 2024). Furthermore, for MLLM-based GUI agents, task scenarios are significantly more complex and dynamic (Shi et al. 2025) compared to traditional MLLM tasks, making OOD detection even more difficult.
Investigating the Challenge of OOD Dection for GUI Agents
@@
The objective of OOD detection for GUI agents, therefore, is to identify whether the pair $(s_{t},x)$ at each time step $t$ deviates from the distribution of $\mathcal{D}_{\text{ID}}$. If OOD is detected, the agent should immediately terminate the execution of the instruction and alert the user; otherwise, it continues with the execution. To formally define the OOD detection mechanism, we introduce the following OOD detection function $f_{\text{OOD}}$, which operates on each pair $(s_{t},x)$.
If $(s_{t},x)\text{ is OOD}$, $f_{\text{OOD}}(s_{t},x)$ returns a value of 1, indicating that the current state are OOD. Otherwise, the agent will predict the appropriate action $a_{t}$ and proceed to execute it. The key to OOD detection for GUI agents is how to construct an effective $f_{\text{OOD}}(s_{t},x)$.
@@
The existing popular OOD detection methods (Malinin and Gales 2018; Yang et al. 2024) can be broadly categorized into two main types: embedding-based methods (Lee et al. 2018) and uncertainty-based methods (Gal and Ghahramani 2016). Based on existing research (Dong et al. 2024; Lu et al. 2024) in OOD detection for (M)LLMs, we used three embedding-based methods and two uncertainty-based methods for experimentation.
We evaluated the five OOD detection methods used as baselines across eight datasets spanning three platforms. Following standard OOD detection method evaluation criteria (Luan et al. 2021; Cui and Wang 2022), we plotted the ROC curve and reported the area under the receiver operating characteristic curve (AUROC) (Metz 1978) and false positive rate at 95% true positive rate (FPR95).
@@
These OOD detection methods rely on finding a decision boundary by identifying the Youden Index (Fluss, Faraggi, and Reiser 2005) after obtaining some scoring metric, that is $\text{Youden Index}=\arg\max_{t}\left(\text{TPR}(t)-\text{FPR}(t)\right)$, where $t$ is the threshold, and TPR and FPR represent the true positive rate and false positive rate.
@@
However, popular GUI agents (Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) have access to diverse training data. In the embedding space, the ID dataset representations for these GUI agents naturally form clusters according to their different data sources. Moreover, we observe that even for datasets originating from a single data source such as AITZ (Zhang et al. 2024b), as shown in Figure 3, there can still be noticeable clustering phenomena. We also observe that samples farther from the centroid tend to yield higher success rates (see Appendix). For such non-linearly separable data, the Youden Index approach becomes inadequate.
@@
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Method Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$ Acc.(%$\uparrow$ Prec.(%$\uparrow$ Rec.(%$\uparrow$ F1(%$\uparrow$
@@
Accuracy(%$\uparrow$ F1 Score(%$\uparrow$ Accuracy(%$\uparrow$ F1 Score(%$\uparrow$ Accuracy(%$\uparrow$ F1 Score(%$\uparrow$
@@
The optimal number of components $m^{*}$ is thus selected as $m^{*}=\arg\min_{m}\text{BIC}(m)$. Once the GMM is fitted with $m^{*}$ components, each Gaussian component $\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ provides a center $\mu_{j}$ and a standard deviation $\sigma_{j}$. We define the ID boundary as the interval $[\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}]$.
@@
For the ID dataset, we use the AITZ (Zhang et al. 2024b) dataset, which contains GUI agent data covering more than 70 Android app scenarios. For the OOD datasets, we select eight GUI agent datasets spanning three platforms: smartphone, computer, and web browser. The smartphone platform includes AndroidControl (Li et al. 2024), OS-Kairos (Cheng et al. 2025b), Meta-GUI (Sun et al. 2022), and ScreenSpot-Mobile (Li et al. 2025),while the computer platform includes Omniact-Desktop (Kapoor et al. 2024) and ScreenSpot-Desktop, and the web browser platform includes Omniact-Web and ScreenSpot-Web. The four smartphone datasets simulate extrapolation-OOD scenarios, whereas the computer and web browser datasets simulate internalization-OOD scenarios.
@@
Mode SCROLL(%$\uparrow$ PRESS(%$\uparrow$ STOP(%$\uparrow$ CLICK(%$\uparrow$ TYPE(%$\uparrow$ TOTAL(%$\uparrow$
@@
Popular GUI Agents (Wu et al. 2024; Qin et al. 2025; Xu et al. 2024) are developed based on Qwen2-VL-7B. To simulate the construction process of a popular GUI Agent, we perform SFT on Qwen2-VL-7B using AITZ train dataset, and then evaluate its OOD detection performance on each sample from eight different OOD datasets and AITZ test dataset (ID). We report the accuracy, precision, recall, and F1 score for each dataset.
Baseline.
We use two commonly-used types of OOD detection methods as our experimental baselines: embedding-based methods and uncertainty-based methods. The embedding-based methods include the TV score (Wang et al. 2024b), last layer embedding, and best layer embedding. The uncertainty-based methods include top-k confidence and output entropy.
@@
In this section, we present several interesting observations and propose a potential method for using GEM to enhance end-to-end tasks in GUI agents. We also demonstrate the rationality of GEM through ablation studies (see Appendix).
@@
To demonstrate the generalization ability of GEM, we evaluate its performance across nine different GUI agents or MLLMs. As shown in Table 3, GEM consistently achieves high accuracy across five models spanning three platforms. For the other four models where GEM performs suboptimally, the results remain explainable. OS-Atlas-Base-7B and OS-Atlas-Pro-7B were exposed to most of the smartphone datasets used in our experiment during their GUI grounding pretraining phase, and their visual encoders were not frozen, which contributed to their weaker OOD performance on the smartphone platform. Meanwhile, LLaVA-1.5 and BLIP capture visual information at a coarser granularity. This highlights their limitations in handling the complex embedding space specific to GUI agent tasks.
@@
We evaluated the effectiveness of representations extracted from each of the twenty-eight layers of Qwen2-VL-7B for OOD detection using an embedding-based approach. As shown in Figure 4, in most datasets, the AUROC first increases as the layer depth increases, then decreases. However, the AUROC rises again in the last two layers. Interestingly, in 4 out of 8 datasets, the best OOD detection results are achieved at the ninth layer. This might be because, as the layers get deeper, the importance of specific task-related features increases, while the contribution of general visual or textual features decreases. Around the ninth layer, the GUI agent likely finds a best balance.

### v4

@@ Abstract
Graphical user interface (GUI) agents have recently emerged as an intriguing paradigm for human-computer interaction, capable of automatically executing user instructions to operate intelligent terminal devices. However, when encountering out-of-distribution (OOD) instructions that violate environmental constraints or exceed the current capabilities of agents, GUI agents may suffer task breakdowns or even pose security threats. Therefore, effective OOD detection for GUI agents is essential. Traditional OOD detection methods perform suboptimally in this domain due to the complex embedding space and evolving GUI environments. In this work, we observe that the in-distribution input semantic space of GUI agents exhibits a clustering pattern with respect to the distance from the centroid. Based on the finding, we propose GEM, a novel method based on fitting a Gaussian mixture model over input embedding distances extracted from the GUI agent that reflect its capability boundary. Evaluated on 8 datasets spanning smartphones, computers, and web browsers, our method achieves an average accuracy improvement of 23.70% over the best-performing baseline while only increasing training time by 4.9% and testing time by 6.5%. We also experimentally demonstrate that GEM can improve the step-wise success rate by 9.40% by requesting assistance from the cloud model when encountering OOD samples. Analysis verifies the generalization ability of our method through experiments on nine different backbones.
Codehttps://github.com/Wuzheng02/GEM-OODforGUIagents
@@ Introduction
Recently, graphical user interface (GUI) agents (zhang2024large; tang2025surveymllmbasedguiagents) have emerged as an intriguing paradigm to human-computer interaction, capable of autonomously executing user instructions and performing human-like control on intelligent terminal devices such as smartphones, computers, and web browsers. The common approach to building GUI agents involves post-training multimodal large language models (MLLMs) using high-quality trajectory data to enhance key task capabilities such as perception (li2023blip; wang2023image), reasoning (yao2022react; wei2022chain), and reflection (liu2025infiguiagent; hu2025guiding). Despite notable advances in instruction following, GUI agents remain vulnerable to out-of-distribution (OOD) risks—executing instructions that violate environmental constraints (e.g., non-existent functions or applications) or exceed the agent’s current capabilities. These failures can lead to task breakdowns or even pose security threats. In real-world applications, OOD risks for GUI agents come with two primary forms: (i) Internalization-OOD, where the agent operates in domain-specific environments (e.g., a particular type of smartphones or vehicle cabins) but incorrectly assumes the presence of unsupported capabilities; (ii) Extrapolation-OOD, where the agent encounters instructions tied to dynamic or evolving environments (e.g., new third-party applications).
As illustrated in Figure 1, following OOD instructions and executing an incorrect action path (e.g., unintentionally resetting a smartphone) can result in critical failures. Therefore, it is essential for GUI agents to incorporate OOD detection mechanisms that identify tasks beyond their supported scope. This not only mitigates potential operational risks but also enables targeted enhancements to agent capabilities through continued research and development.
Popular OOD detection methods can be broadly categorized (malinin2018predictive) into two types: embedding-based approaches (lee2018simple) and model uncertainty-based approaches (hendrycks2016baseline). These methods have demonstrated effectiveness in traditional (M)LLM tasks such as summarization (ren2022out), visual question answering (kervadec2021roses), mathematical reasoning (wang2024embedding), and text generation (wu2022multi). Directly applying these existing OOD detection techniques to the GUI agent domain results in suboptimal performance. Notably, even with the classification threshold calibrated to correctly identify 95% of in-distribution (ID) tasks, the best-performing method still misclassifies between 31.75% and 48.92% of OOD tasks as ID.
@@
(i) Complex Embedding Space: The inputs to GUI agents are inherently complex (ma2024comprehensive). Concretely, GUI screens typically contain densely populated UI components (zhang2024large), leading to higher information density than traditional MLLM tasks. Besides, the diversity of user instructions further complicates the inference of user intent. This complexity substantially increases the difficulty of effective OOD detection for GUI agents.
(ii) Evolving Environments: GUI environments frequently change due to system upgrades or the installation of new third-party applications. As a result, GUI agents must continually adapt their capabilities (wang2025mobileE; liu2025learnact). This evolving nature complicates OOD detection by demanding both accurate capability assessment and temporal adaptability. Consequently, reliance on static external classifiers becomes increasingly impractical.
To address the challenges above, we propose GEM, a novel method based on fitting a Gaussian mixture model (GMM) (reynolds2009gaussian) over input embedding distances extracted from the GUI Agent that reflect its capability boundary. Specifically, we first extract the input embeddings from the encoder layer of the GUI agent on its training data, and fit them into a high-dimensional hypersphere. We then compute the L2-norm distances of all embeddings relative to the centroid of the hypersphere. Under bayesian information criterion (BIC) (neath2012bayesian) supervision, we fit a GMM to these distances and define the OOD detection boundary as a configurable number of standard deviations away from each GMM cluster center. Experiments across eight GUI agent datasets spanning smartphones, computers, and web browsers platforms show that our method consistently outperforms all traditional OOD detection methods. And we provide a possible method to use GEM to improve the end-to-end results for GUI agents.
@@
Method Smartphone Computer Web browser
AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$
@@
With the advancement of MLLMs (liu2023visual; achiam2023gpt), numerous GUI agent foundation models (hong2024cogagent; wu2024atlas; qin2025ui) have emerged, capable of executing user instructions across smartphones, computers, and web browsers. For the technical framework, prompt-based approaches (wang2025mobileE; jiang2025appagentx; wu2025quick) have been developed, leveraging closed-source models to construct agent systems that fulfill user instructions. Additionally, researchers have also explored pre-training (ma2024comprehensive; wu2025smoothing), supervised fine-tuning (SFT) (zhang2023you; wu2025verios), and reinforcement learning (RL) (zhoudigirl; xia2025gui; tang2025guig2gaussianrewardmodeling) techniques to further enhance the ability of GUI agents to complete user instructions. Regardless of their advancements, GUI agents inevitably face capability limitations. In complex and dynamic real-world environments (chengkairos; guo2025atomic), GUI agents are prone to encountering OOD situations.
@@
For OOD detection in MLLMs, it is necessary to jointly consider both visual and textual modalities, in contrast to traditional computer vision OOD detection, which only requires modeling the visual modality (liu2020energy; ren2019likelihood), or LLMs OOD detection, which focuses on textual modality (wang2024embedding; ren2022out). In the context of MLLMs, some researchers have proposed using maximum concept matching (ming2022delving) to characterize OOD uncertainty, while others have approached the problem by training models on ID datasets and classifying whether an input image belongs to an unknown category (wang2023clipn). However, OOD detection for MLLMs remains a challenging research area (dong2024multiood; lu2024recent). Furthermore, for MLLM-based GUI agents, task scenarios are significantly more complex and dynamic (shi2025towards) compared to traditional MLLM tasks, making OOD detection even more difficult.
Investigating the Challenge of OOD Detection for GUI Agents
@@
The objective of OOD detection for GUI agents, therefore, is to identify whether the pair $(s_{t},x)$ at each time step $t$ deviates from the distribution of $\mathcal{D}_{\text{ID}}$. If OOD is detected, the agent should immediately terminate the execution of the instruction and alert the user; otherwise, it continues with the execution. To formally define the OOD detection mechanism, we introduce the following OOD detection function $f_{\text{OOD}}$, which operates on each pair $(s_{t},x)$. If $(s_{t},x)\text{ is OOD}$, $f_{\text{OOD}}(s_{t},x)$ returns a value of 1. Otherwise, the agent will predict the appropriate action $a_{t}$ and proceed to execute it.
@@
Existing popular OOD detection methods (malinin2018predictive; yang2024generalized) can be broadly categorized into two main types: embedding-based methods (lee2018simple) and uncertainty-based methods (gal2016dropout). Based on existing research (dong2024multiood; lu2024recent) in OOD detection for (M)LLMs, we used three embedding-based methods and two uncertainty-based methods for experimentation.
We evaluated the methods used as baselines across eight datasets spanning three platforms. Following standard OOD detection method evaluation criteria (luan2021out; cui2022out), we plotted the ROC curve and reported the area under the receiver operating characteristic curve (AUROC) (metz1978basic) and false positive rate at 95% true positive rate (FPR95).
@@
These OOD detection methods rely on finding a decision boundary by identifying the Youden Index (fluss2005estimation) after obtaining some scoring metric, that is $\text{Youden Index}=\arg\max_{t}\left(\text{TPR}(t)-\text{FPR}(t)\right)$, where $t$ is the threshold, and TPR and FPR represent the true positive rate and false positive rate.
@@
However, popular GUI agents (wu2024atlas; qin2025ui; xu2024aguvis) have access to diverse training data. In the embedding space, the ID dataset representations for these GUI agents naturally form clusters according to their different data sources. Moreover, we observe that even for datasets originating from a single data source such as AITZ (zhang2024android), as shown in Figure 3, there can still be noticeable clustering phenomena. We also observe that samples farther from the centroid tend to yield higher success rates. For such non-linearly separable data, the Youden Index approach becomes inadequate.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The optimal number of components $m^{*}$ is thus selected as $m^{*}=\arg\min_{m}\text{BIC}(m)$. Once the GMM is fitted with $m^{*}$ components, each Gaussian component $\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ provides a center $\mu_{j}$ and a standard deviation $\sigma_{j}$. We define the ID boundary as the interval $\mu_{j}-n\sigma_{j},\mu_{j}+n\sigma_{j}]$.
@@
For the ID dataset, we use the AITZ (zhang2024android) dataset, which contains GUI agent data covering more than 70 Android app scenarios. For the OOD datasets, we select eight GUI agent datasets spanning three platforms: smartphone, computer, and web browser. The smartphone platform includes AndroidControl (li2024effects), OS-Kairos (cheng2025kairos), Meta-GUI (sun2022meta), and ScreenSpot-Mobile (li2025ScreenSpot), while the computer platform includes Omniact-Desktop (kapoor2024omniact) and ScreenSpot-Desktop, and the web browser platform includes Omniact-Web and ScreenSpot-Web.
@@
[…]
@@
Popular GUI Agents (wu2024atlas; qin2025ui; xu2024aguvis) are developed based on Qwen2-VL-7B. To simulate the construction process of a popular GUI Agent, we perform SFT on Qwen2-VL-7B using AITZ train dataset, and then evaluate its OOD detection performance on each sample from OOD datasets and AITZ test dataset (ID).
@@
In this section, we present several interesting observations and propose a potential method for using GEM to enhance end-to-end tasks in GUI agents.
@@
To demonstrate the generalization ability of GEM, we evaluate its performance across nine different GUI agents or MLLMs. As shown in Table 3, GEM consistently achieves high accuracy across five models spanning three platforms. For the results where GEM performs suboptimally, OS-Atlas were exposed to most of the smartphone datasets used in our experiment during their GUI grounding pretraining phase. Meanwhile, LLaVA-1.5 and BLIP capture visual information at a coarser granularity. This highlights their limitations in handling the complex embedding space specific to GUI agent tasks.
@@
We evaluated the effectiveness of representations extracted from each of the twenty-eight layers of Qwen2-VL-7B for OOD detection using an embedding-based approach. As shown in Figure 4, in most datasets, the AUROC first increases as the layer depth increases, then decreases. However, the AUROC rises again in the last two layers. Interestingly, in 4 out of 8 datasets, the best OOD detection results are achieved at the ninth layer. This might be because, as the layers get deeper, the importance of specific task-related features increases, while the contribution of general visual or textual features decreases. Around the ninth layer, the GUI agent likely finds a best balance. And the GUI agent has developed a kind of confidence estimation in its final output in the final two layeers, allowing it to better distinguish between OOD and ID samples.